OpenStudy (anonymous):

the square root of 49/x^3

jimthompson5910 (jim_thompson5910):

\[\Large \sqrt{\frac{49}{x^3}}\] \[\Large \frac{\sqrt{49}}{\sqrt{x^3}}\] \[\Large \frac{\sqrt{7^2}}{\sqrt{x^2*x}}\] \[\Large \frac{\sqrt{7^2}}{\sqrt{x^2}*\sqrt{x}}\] \[\Large \frac{7}{x*\sqrt{x}}\] ------------------------------ So far, I've shown that \[\Large \sqrt{\frac{49}{x^3}}\] turns into \[\Large \frac{7}{x*\sqrt{x}}\] Now let's rationalize the denominator ----------------------------- \[\Large \frac{7}{x*\sqrt{x}}\] \[\Large \frac{7{\color{red}{*\sqrt{x}}}}{x*\sqrt{x}{\color{red}{*\sqrt{x}}}}\] \[\Large \frac{7\sqrt{x}}{x*\sqrt{x*x}}\] \[\Large \frac{7\sqrt{x}}{x*\sqrt{x^2}}\] \[\Large \frac{7\sqrt{x}}{x*x}\] \[\Large \frac{7\sqrt{x}}{x^2}\] ----------------------------- So in the end, \[\Large \sqrt{\frac{49}{x^3}}\] simplifies to \[\Large \frac{7\sqrt{x}}{x^2}\]